# SGFormer: Spherical Geometry Transformer for $360^{\circ}$ Depth Estimation

### v1

@@
Abstract.
Panoramic distortion poses a significant challenge in $360^{\circ}$ depth estimation, particularly pronounced at the north and south poles. Existing methods either adopt a bi-projection fusion strategy to remove distortions or model long-range dependencies to capture global structures, which can result in either unclear structure or insufficient local perception. In this paper, we propose a spherical geometry transformer, named SGFormer, to address the above issues, with an innovative step to integrate spherical geometric priors into vision transformers. To this end, we retarget the transformer decoder to a spherical prior decoder (termed SPDecoder), which endeavors to uphold the integrity of spherical structures during decoding. Concretely, we leverage bipolar re-projection, circular rotation, and curve local embedding to preserve the spherical characteristics of equidistortion, continuity, and surface distance, respectively. Furthermore, we present a query-based global conditional position embedding to compensate for spatial structure at varying resolutions. It not only boosts the global perception of spatial position but also sharpens the depth structure across different patches. Finally, we conduct extensive experiments on popular benchmarks, demonstrating our superiority over state-of-the-art solutions.
@@ 1. Introduction
Figure 1. (a) Bi-projection fusion strategy. Bifuse (Wang et al., 2020) combines feature maps from equirectangular and cubemap projections. (b) Long-range dependency strategy. EGFormer (Yun et al., 2023) adapts vertical and horizontal shaped local windows. (c) Our SGFormer presents a query-based global conditional position embedding(GCPE) and integrates spherical geometric priors, significantly improving the performance of panoramic depth estimation.
@@
The panoramas represent a whole scene as an equirectangular map, delivering a $360^{\circ}$ Field of View (FoV) of the surroundings. Hence, estimating depth from a single panoramic image can offer complete depth of the scene, with a holistic understanding of the surroundings. However, the equidistant unfolding characteristic (also known as panoramic distortion) of panoramas poses a new challenge for $360^{\circ}$ depth estimation. Recently, researchers (Wang et al., 2020; Ai et al., 2023) have preferred to estimate accurate panoramic depth with the assistance of other projection formats (e.g., Cubemap projection (CP) and Tangent projection (TP)). These projections significantly reduce distortions by mapping a panorama onto several tangent faces, but damage the inherent continuous representation of panoramic images. Many approaches (Wang et al., 2020; Jiang et al., 2021; Li et al., 2022; Ai et al., 2023) typically leverage the complementarity of these projection representations (shown in Fig. 1a) to estimate more accurate panoramic depth. Nevertheless, the trade-off between the FoV continuity and distortion elimination poses another challenge for the model design. Besides, these bi-projection fusion strategies require additional branches, increasing the computational burden. In view of this, we only take the equirectangular projection as the input and delve into the properties of the spherical space to deal with the distortion.
@@
On the other hand, some approaches (Pintore et al., 2021; Shen et al., 2022a; Yun et al., 2023) deal with the panoramic distortions via modeling long-range dependencies (as illustrated in Fig. 1b). In previous work (Pintore et al., 2021), long short-term memory (Hochreiter and Schmidhuber, 1997) (LSTM) is introduced to disseminate global information in vertically compressed sequences. However, the geometric structure, which is an essential cue to estimate depth, is neglected, thus delivering blurred depth details. Meanwhile, those methods based on modeling long-range dependencies are either too global to perceive the local details (Yun et al., 2023), or too local to deal with the distortion efficiently (Shen et al., 2022a).
@@
In this paper, we incorporate spherical geometry into our model (i.e., SGFormer, as shown in Fig. 1c) to capture local details while promoting global interaction. We leverage three geometric priors of the sphere to design a customized panoramic decoder (SPDecoder). Firstly, we use the spherical equidistortion to design a bipolar re-projection paradigm to deal with the distortions. We re-project these areas with severe distortions to the equator region to eliminate the negative effect of distortions. Especially, our method is parameter-free and transformation-efficient without any other branches. Secondly, we perform circular rotation through the continuity of the sphere. It helps to produce seamless panoramic representations, which allows two separate points, originally located on the left and right boundaries, to be close to each other and have similar semantics. Thirdly, with the spherical surface distance, curve local embedding is presented to generate curved relative position embedding for each point in the spherical projection space, which provides additional geometric distance and enhances the understanding of spatial structure.
@@
Moreover, to further enhance performance comprehensively, we design a query-based global conditional position embedding (GCPE) scheme. Using features of different resolutions as queries to query the keys that match the resolution, we can adaptively provide different geometric details and spatial structures at different resolutions. This additional structure, in turn, explicitly guides our SPDecoder with global spatial positioning, thereby improving the accuracy of estimated depth.
@@
Finally, we evaluate the proposed SGFormer on two benchmarks, Structured3D (Zheng et al., 2020) and Pano3D (Albanis et al., 2021). Experimental results demonstrate that our method outperforms existing methods by a large margin on both quantitative and qualitative results. Our main contributions can be summarized as follows:
@@
2. Related Work
2.1. Monocular $360^{\circ}$ Depth Estimation
@@
Bi-projection fusion strategy. BiFuse (Wang et al., 2020) introduces a dual-branch fusion method that combines CP and ERP features to alleviate distortion, while UniFuse (Jiang et al., 2021) proposes a new framework that unidirectionally fuses CP features into ERP features during the decoding stage, enhancing both efficiency and performance. Glpanodepth (Bai et al., 2022) proposes a fusion module that combines ERP and CP features for panoramic depth estimation. 360MonoDepth (Rey-Area et al., 2022) introduces a framework that uses tangent images for depth estimation from high-resolution $360^{\circ}$ images. OmniFusion (Li et al., 2022) directly converts $360^{\circ}$ images into less-distorted perspective patches and uses geometric priors to merge the predictions of these patches into the ERP space to produce the final depth output in ERP format. HRDFuse (Ai et al., 2023) employs a collaborative mechanism that explores both the overall contextual information of the ERP and the local structure of the TP, estimating panoramic depth by learning the similarity of features between the ERP and TP.
@@
Modeling long-range dependencies. SliceNet (Pintore et al., 2021) divides the ERP image into vertical slices, however, it is hard to recover the details near the poles because it ignores the latitudinal distortion property. PanoFormer (Shen et al., 2022a) attempts to estimate panoramic depth by leveraging a transformer-based architecture to process TP patches. By sampling the eight most relevant tokens for each central token on the ERP domain and adaptively adjusting local window sizes with a learnable offset, it enhances the precision in capturing the 2D structure and spatial details of each patch. EGFormer (Yun et al., 2023) leverages equirectangular geometry as the bias for local windows, thus estimating panoramic depth through an equirectangular geometry-aware local attention with a large receptive field. However, the vertical and horizontal shaped local windows also weaken the perception of the structure of each local area.
@@
Other perspectives. Some researchers (Coors et al., 2018; Dai et al., 2017; Eigen et al., 2014; Zhu et al., 2020; Xiong and Grauman, 2018; Jiang et al., 2019; Khasanova and Frossard, 2019; Laina et al., 2016) employ various distortion-aware convolution filters to mitigate distortion, but due to the limited receptive fields, the final effect is still not entirely satisfactory. In deeper levels (Su and Grauman, 2019). Several spherical CNNs (Cohen et al., 2018; Yang et al., 2020) have employed convolution within the spectral domain, which could lead to a heavier computational burden. (Shen et al., 2022b) considers the significance of the geometric structure in depth estimation and proposes a neural contourlet network for panoramic monocular depth estimation.
2.2. Transformer for Vision Tasks
@@
Figure 2. Overview of our proposed SGFormer. The SGFormer is composed of three major parts: feature extractors for ERP images at varying resolutions, calculation of global conditional position embedding (CGCPE) module, and the SPDecoder that integrates three spherical geometric priors.
@@
Inspired by the transformer in the NLP field, the Vision Transformer (Dosovitskiy et al., 2020) (ViT) uses a global self-attention mechanism for visual tasks. However, the global information interaction often results in higher computational costs. Swin Transformer (Liu et al., 2021) (SwinT) adopts a window-based self-attention, incorporating a hierarchical architecture and shifting window mechanism for image processing. CSwin Transformer (Dong et al., 2022) (CSwinT) introduces a cross-shaped window attention mechanism, conducting parallel self-attention computations on horizontal and vertical stripes. (Ling et al., 2023; Cho et al., 2022) propose distortion relief methods specifically designed for panoramic image vision tasks. Despite these advancements, the distortion in panoramic images poses some challenges for transformers in panoramic image processing.
@@
PanoFormer (Shen et al., 2022a) redesigns the traditional transformer with tangent patches on the spherical domain and panorama specific metrics, using learnable token flows to perceive geometric structures adaptively. HRDFuse (Ai et al., 2023) processes panoramic images into TP patches and aggregates global information in a ViT manner. EGFormer (Yun et al., 2023) combines with CSwinT (Dong et al., 2022)to propose an equirectangular geometry-biased transformer. However, this horizontal and vertical self-attention weakens the perception of local structures. In this work, we use the spherical geometric characteristics to further explore the suitable transformer structure for panoramic views.
3. SGFormer
3.1. Architecture Overview
@@
Figure 3. Overview of the query-based CGCPE module. f3 generates a global key (i.e. GCPE3) by passing through a transformer module with global spherical position embedding (GSPE). The global key is then processed through linear layers to generate keys corresponding to features at various resolutions. Features of different resolutions are used as queries, which are subsequently operated by $QK^{T}$. After another linear transformation, the corresponding GCPEs at different resolutions are obtained.
@@
In Fig. 2, we show our complete framework motivated by the spherical properties. Overall, the proposed SGFormer takes a panorama as input and outputs a corresponding high-quality depth map. During the encoder stage, we employ Resnet-34 to extract features at five different resolutions (Fig. 2 ①). Then these features are fed into the CGCPE module (Fig. 2 ② ), which is a query-based module that adaptively produces GCPEs, compensating for additional spatial geometric structure (More details are in Section 3.2). The visualization results of GCPEs demonstrate that they provide geometric details, promoting a comprehensive understanding of spatial location. Subsequently, a hierarchical decoder (the SPDecoder, described in Section 3.3 ) is designed to estimate depth from these extracted features and integrates three spherical geometric priors to maintain the unity of the spherical geometric structure. Due to explicit structural cues and spherical priors (i.e., equidistortion, continuity, and surface distance), our scheme can relieve distortion effectively and these structure- and sphere-aware designs can promote the performance of our model.
3.2. CGCPE Module
@@
Previous work (Chen et al., 2021) has proved that the structure of the scenes is crucial for depth estimation. Hence, we aim to leverage the panoramic geometric cue to assist in the depth estimation process. To achieve this, we introduce GCPEs, which are learned from the extracted features and a spherical surface distance (i.e. Fig. 3 GSPE), to conduct the depth estimation with an additional explicit structure cue. The introduced GCPEs provide scale-adaptive variable details, which can push the network to explore the diverse spatial structures of the same panoramic space but with different resolutions.
@@
Specifically, as shown in Fig. 3, we select the feature with a $\frac{H}{16}\times\frac{W}{16}$ resolution (denoted as $f_{3}$) as the reference feature. There is a trade-off between computational efficiency and clear details. To balance them, we choose the middle scale. The selected features are fed into a transformer module and then we achieve position embedding via a GSPE approach (more details are in section 3.3, curve local embedding), thereby promoting the generation of panoramic structure (i.e., global key in Fig. 3). As shown in the top-left corner of Fig. 3, GSPE projects any two pixels in the ERP feature map into the spherical space to calculate the spherical surface distance as a sphere-aware embedding. Subsequently, the aggregated global geometric information is transformed into keys via a linear projection layer. At each scale, we leverage features as queries to query the keys, adaptively generating GCPE at each resolution. From Fig. 4, the depth details can be observed clearly, indicating that the query-based GCPE strategy can contribute to the spatial structure and depth details.
@@
Figure 4. The visualization of heatmap of features and the predicted depth map with/without GCPE.
@@
Figure 5. Overview of SPDecoder (left), which is combined with three spherical geometric priors, and SPAttention (right).
3.3. SPDecoder
@@
To balance the global perception and the local details enhancement, we propose the SPDecoder by leveraging the spherical priors. As illustrated in Fig. 5, we send the input features to SPAttention firstly, where we add curve local embedding to provide additional sphere-aware geometric distance. Then, considering the sphere’s continuity, we perform circular rotation to enhance the feature interaction among patches (Fig. 5 ①). We restore the original state through reverse circular rotation (Fig. 5 ②). Afterward, a bipolar re-projection approach based on the spherical equidistortion property is designed to relieve the distortion (Fig. 5 ③). Moreover, circular rotation is operated again to interact with the original features in the vertical direction (Fig. 5 ④). Ultimately, we revert to the original state through reverse operations (Fig. 5 ⑤⑥). We describe these operations in detail as follows.
@@
Figure 6. (a) A projection imaging model with the x-axis (North) as the positive direction. (b) A projection imaging model with the y-axis (North) as the positive direction.
@@
(1) $\left\{\begin{matrix}\lambda^{\prime}=\frac{v^{\prime}}{H}\times\pi\\ \varphi^{\prime}=\frac{u^{\prime}}{W}\times 2\pi\end{matrix}\right.$
@@
(2) $\left\{\begin{matrix}x=\sin\left(\lambda^{\prime}\right)\\ y=\cos\left(\lambda^{\prime}\right)\\ z=\cos\left(\frac{\pi}{2}-\lambda^{\prime}\right)\cdot\sin\left(\varphi^{\prime}\right)\end{matrix}\right.$
@@
(3) $\left\{\begin{matrix}\lambda=\arccos\left(\frac{x}{\sqrt{z^{2}+y^{2}}}\right)\\ \varphi=\arcsin\left(\frac{y}{\sqrt{z^{2}+y^{2}}}\right)\end{matrix}\right.$
@@
(4) $\left(u,v\right)=\left(\frac{\lambda}{\pi}\times H,\frac{\varphi}{2\pi}\times W\right)$
In particular, we employ the bilinear interpolation function to fill the missing pixels. After BRP operation, the polar regions are projected to the equator (the first column of Fig. 7), which allows the model to estimate accurate depths with these distortion-free features. Besides, the proposed BRP approach is parameter-free and efficient in transformation, and no additional branches are required.
Figure 7. The first row shows the division of patches before and after CR. The second row displays the division of patches after BRP followed by another CR. The blue and red lines represent the divisions before and after CR, respectively.
Circular rotation (CR): The traditional transformer (Vaswani et al., 2017) architecture primarily relies on a global self-attention mechanism, which is inefficient when dealing with high-resolution images. To address this, SwinT (Liu et al., 2021) performs self-attention among divided window-based patches and employs a shifting window operation to facilitate feature interaction between the adjacent patches. Due to the unique seamless property of panoramic images, we implement self-attention interactions between local windows by rotating half the window size in the horizontal direction (the first row in Fig. 7). Then we rotate back to the original state to align with the initial spatial representation. Since the north and south poles of panoramic images are not adjacent in space, we cannot directly apply CR in the vertical direction. However, benefiting from the BRP module, we can map the features from the vertical to the horizontal direction. Then, through CR, features that are originally in the vertical direction are interacted. Finally, we obtain a feature representation that combines horizontal and vertical directions.
Figure 8. The left image shows an ERP format panoramic image, with each square indicating a segmented patch. Red and blue squares mark patches from the top and bottom with high distortion, while yellow and green squares indicate middle area patches with less distortion. These patches are projected onto the spherical space on the right.
@@
For each patch containing N points, we define $C$ as the CLE within each ${N\times N}$ patch. $\forall P_{1}\left(\lambda_{1},\varphi_{1}\right)$ and $P_{2}\left(\lambda_{2},\varphi_{2}\right)$ (i.e., the position of two points within a patch in spherical coordinates) to calculate the CLE. $c\left(P_{1},P_{2}\right)$ represents the distance between $P_{1}$ and $P_{2}$, where $\lambda\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$/$\varphi\in\left(-\pi,\pi\right)$ represents latitude/longitude, repectively. Through the Haversine formula (Sinnott, 1984), we can calculate $c\left(P_{1},P_{2}\right)$ as follows:
(5) $h=\sin^{2}\left(\frac{\lambda_{2}-\lambda_{1}}{2}\right)+\cos\left(\lambda_{1}\right)\cdot\cos\left(\lambda_{2}\right)\cdot\sin^{2}\left(\frac{\varphi_{2}-\varphi_{1}}{2}\right)$
(6) $c\left(P_{1},P_{2}\right)=2R\cdot\arctan\left(\sqrt{\frac{h}{1-h}}\right)$
@@
Table 1. Quantitative comparison on the S3D (Structured3D) and Pano3D datasets with the SoTA methods. Blue indicates that our method achieves the best performance.
@@
Bifuse (Wang et al., 2020) 0.0644 0.0565 0.4099 0.1194 0.9673 0.9892 0.9948
SliceNet (Pintore et al., 2021) 0.1103 0.1273 0.6164 0.1811 0.9012 0.9705 0.9867
Yun et al. (Yun et al., 2022) 0.0505 0.0499 0.3475 0.1150 0.9700 0.9896 0.9947
S3D (Zheng et al., 2020) PanoFormer (Shen et al., 2022a) 0.0394 0.0346 0.2960 0.1004 0.9781 0.9918 0.9958
EGFormer (Yun et al., 2023) 0.0342 0.0279 0.2756 0.0932 0.9810 0.9928 0.9964

### v2

# SGFormer: Spherical Geometry Transformer for 360 Depth Estimation

@@
Abstract
Panoramic distortion poses a significant challenge in 360 depth estimation, particularly pronounced at the north and south poles. Existing methods either adopt a bi-projection fusion strategy to remove distortions or model long-range dependencies to capture global structures, resulting in either unclear structure or insufficient local perception. In this paper, we propose a spherical geometry transformer, named SGFormer, to address the above issues, with an innovative step to integrate spherical geometric priors into vision transformers. To this end, we retarget the transformer decoder to a spherical prior decoder (termed SPDecoder), which endeavors to uphold the integrity of spherical structures during decoding. Concretely, we leverage bipolar reprojection, circular rotation, and curve local embedding to preserve the spherical characteristics of equidistortion, continuity, and surface distance, respectively. Furthermore, we present a query-based global conditional position embedding to compensate for spatial structure at varying resolutions. It not only boosts the global perception of spatial position but also sharpens the depth structure across different patches. Finally, we conduct extensive experiments on popular benchmarks, demonstrating our superiority over state-of-the-art solutions. Our code will be made publicly at https://github.com/iuiuJaon/SGFormer.
Index Terms:
@@ I Introduction
Figure 1: Brief comparisons between previous typical methods and the proposed method: (a) Bi-projection fusion strategy. Bifuse [1] combines feature maps from equirectangular and cubemap projections. (b) Long-range dependency strategy. EGFormer [2] adapts vertical and horizontal shaped local windows. (c) Our SGFormer presents a query-based global conditional position embedding (GCPE) and integrates spherical geometric priors, significantly improving the performance of panoramic depth estimation.
@@
The panoramas represent a whole scene as an equirectangular map, delivering a 360 Field of View (FoV) of the surroundings. Hence, estimating depth from a single panoramic image can offer complete depth of the scene, with a holistic understanding of the surroundings. However, the equidistant unfolding characteristic (also known as panoramic distortion) of panoramas poses a new challenge for 360 depth estimation. Therefore, directly applying the perspective-based approaches [3, 4, 5, 6, 7, 8] to estimate depth from a single panorama is challenging. To address this issue, researchers [1, 9] have recently preferred to estimate accurate panoramic depth with the assistance of other projection formats (e.g., Cubemap projection (CP), and Tangent projection (TP)). These projections significantly reduce distortions by mapping a panorama onto several tangent faces, but damage the inherent continuous representation of panoramic images. Many approaches [1, 10, 11, 9] typically leverage the complementarity of these projection representations (shown in Fig. 1a) to estimate more accurate panoramic depth. Nevertheless, the trade-off between the FoV continuity and distortion elimination poses another challenge for the model design. Besides, these bi-projection fusion strategies require additional branches, increasing the computational burden. In view of this, we only take the equirectangular projection as the input and delve into the properties of the spherical space to deal with the distortion.
@@
On the other hand, some approaches [12, 13, 2] deal with the panoramic distortions via modeling long-range dependencies (as illustrated in Fig. 1b). In previous work [12], long short-term memory [14] (LSTM) is introduced to disseminate global information in vertically compressed sequences. However, the geometric structure, which is an essential cue to estimate depth, is neglected, thus delivering blurred depth d. Meanwhile, those methods based on modeling long-range dependencies are either too global to perceive the local details [2], or too local to deal with the distortion efficiently [13].
@@
In this paper, we incorporate spherical geometry into our model (i.e., SGFormer, as shown in Fig. 1c) to capture local details while promoting global interaction. We leverage three geometric priors of the sphere to design a customized panoramic decoder (SPDecoder). First, we use the spherical equidistortion to design a bipolar re-projection paradigm to deal with the distortions. We re-project these areas with severe distortions to the equator region to eliminate the negative effect of distortions. Specifically, our method is parameter-free and transformation-efficient without any other branches. Subsequently, we perform circular rotation through the continuity of the sphere. It helps to produce seamless panoramic representations, which allows two separate points, originally located on the left and right boundaries, to be close to each other and have similar semantics. Furthermore, with the spherical surface distance, curve local embedding is presented to generate curved relative position embedding for each point in the spherical projection space, which provides additional geometric distance and enhances the understanding of spatial structure.
@@
To enhance performance comprehensively, we further design a query-based global conditional position embedding (GCPE) scheme. By leveraging features of different resolutions as queries to query the keys that match the resolution, we can adaptively provide different geometric details and spatial structures at different resolutions. This additional structure, in turn, explicitly guides our SPDecoder with global spatial positioning, thereby improving the accuracy of estimated depth.
@@
We evaluate the proposed SGFormer on two benchmarks, Structured3D [15] and Pano3D [16]. Experimental results demonstrate that our method outperforms existing methods by a large margin on both quantitative and qualitative results. Our main contributions can be summarized as follows:
@@
II Related Work
II-A Monocular 360 Depth Estimation
@@
Bi-projection fusion strategy. BiFuse [1] introduces a dual-branch fusion method that combines CP and ERP features to alleviate distortion, while UniFuse [10] proposes a new framework that unidirectionally fuses CP features into ERP features during the decoding stage, enhancing both efficiency and performance. Glpanodepth [17] proposes a fusion module that combines ERP and CP features for panoramic depth estimation. 360MonoDepth [18] introduces a framework that uses tangent images for depth estimation from high-resolution 360 images. OmniFusion [11] directly converts 360 images into less-distorted perspective patches and uses geometric priors to merge the predictions of these patches into the ERP space to produce the final depth output in ERP format. HRDFuse [9] employs a collaborative mechanism that explores both the overall contextual information of the ERP and the local structure of the TP, estimating panoramic depth by learning the similarity of features between the ERP and TP.
@@
Modeling long-range dependencies. SliceNet [12] divides the ERP image into vertical slices, however, it is hard to recover the details near the poles because it ignores the latitudinal distortion property. PanoFormer [13] attempts to estimate panoramic depth by leveraging a transformer-based architecture to process TP patches. By sampling the eight most relevant tokens for each central token on the ERP domain and adaptively adjusting local window sizes with a learnable offset, it enhances the precision in capturing the 2D structure and spatial details of each patch. EGFormer [2] leverages equirectangular geometry as the bias for local windows, thus estimating panoramic depth through an equirectangular geometry-aware local attention with a large receptive field. However, the vertical and horizontal shaped local windows also weaken the perception of the structure of each local area.
@@
Other perspectives. Some researchers [19, 20, 21, 22, 23, 24, 25] employ various distortion-aware convolution filters to mitigate distortion, but due to the limited receptive fields, the final effect is still not entirely satisfactory. In deeper levels [26]. Several spherical CNNs [27] have employed convolution within the spectral domain, which could lead to a heavier computational burden. [28] considers the significance of the geometric structure in depth estimation and proposes a neural contourlet network for panoramic monocular depth estimation.
II-B Transformer for Vision Tasks
@@
Figure 2: Overview of our proposed SGFormer. The SGFormer comprises three major parts: feature extractors for ERP images at varying resolutions, calculation of global conditional position embedding (CGCPE) module, and the SPDecoder that integrates three spherical geometric priors.
@@
Inspired by the transformer in the NLP field, the Vision Transformer [29] (ViT) uses a global self-attention mechanism for visual tasks. However, the global information interaction often results in higher computational costs. Swin Transformer [30] (SwinT) adopts a window-based self-attention, incorporating a hierarchical architecture and shifting window mechanism for image processing. CSwin Transformer [31] (CSwinT) introduces a cross-shaped window attention mechanism, conducting parallel self-attention computations on horizontal and vertical stripes. [32, 33] propose distortion relief methods specifically designed for panoramic image vision tasks. Despite these advancements, the distortion in panoramic images poses some challenges for transformers in panoramic image processing.
@@
PanoFormer [13] redesigns the traditional transformer with tangent patches on the spherical domain and panorama-specific metrics, using learnable token flows to perceive geometric structures adaptively. HRDFuse [9] processes panoramic images into TP patches and aggregates global information in a ViT manner. EGFormer [2] combines with CSwinT [31] to propose an equirectangular geometry-biased transformer. However, this horizontal and vertical self-attention weakens the perception of local structures. In this work, we use the spherical geometric characteristics to further explore the suitable transformer structure for panoramic views.
III SGFormer
III-A Architecture Overview
@@
Figure 3: Overview of the query-based CGCPE module. In particular, the extracted features at the 3rd scale (denoted as f3) generate a global key (i.e. GCPE3) by passing through a transformer module with global spherical position embedding (GSPE). The global key is then processed through linear layers to generate keys corresponding to features at various resolutions. Features of different resolutions are used as queries, which are subsequently operated by $QK^{T}$. After another linear transformation, the corresponding GCPEs at different resolutions are obtained.
@@
In Fig. 2, we show our complete framework motivated by the spherical properties. Overall, the proposed SGFormer takes a panorama as input and outputs a corresponding high-quality depth map. During the encoder stage, we employ Resnet-34 to extract features at five different resolutions (Fig. 2 ①). Then, these features are fed into the calculation of global conditional position embedding (CGCPE) module (Fig. 2 ② ), which is a query-based module that adaptively produces global conditional position embedding (GCPEs), compensating for additional spatial geometric structure (More details are presented in Section III-B). The visualization results of GCPEs demonstrate that they provide geometric details, promoting a comprehensive understanding of spatial location. Subsequently, a hierarchical decoder (the SPDecoder, described in Section III-C) is designed to estimate depth from these extracted features and integrates three spherical geometric priors to maintain the unity of the spherical geometric structure. As benefits by explicit structural cues and spherical priors (i.e. equidistortion, continuity, and surface distance), our scheme can effectively relieve distortion and these structure- and sphere-aware designs can promote the performance of our model.
III-B Calculation of Global Conditional Position Embedding (CGCPE) Module
@@
We find that the structure of the scenes is crucial for depth estimation. Hence, we aim to leverage the panoramic geometric cue to assist in the depth estimation process. To achieve this, we introduce GCPEs, which are learned from the extracted features and a spherical surface distance (i.e. Fig. 3 global spherical position embedding (GSPE)), to conduct the depth estimation with an additional explicit structure cue. The introduced GCPEs provide scale-adaptive variable details, which can push the network to explore the diverse spatial structures of the same panoramic space but with different resolutions.
@@
Specifically, as shown in Fig. 3, we select the feature with a $\frac{H}{16}\times\frac{W}{16}$ resolution (denoted as $f_{3}$) as the reference feature. There is a trade-off between computational efficiency and clear details. To balance them, we choose the middle scale. The selected features are fed into a transformer module, and then we achieve position embedding via a GSPE approach (more details are in section III-C, curve local embedding), thereby promoting the generation of panoramic structure (i.e., global key in Fig. 3). As shown in the top-left corner of Fig. 3, GSPE projects any two pixels in the ERP feature map into the spherical space to calculate the spherical surface distance as a sphere-aware embedding. Subsequently, the aggregated global geometric information is transformed into keys via a linear projection layer. At each scale, we leverage features as queries to query the keys, adaptively generating GCPE at each resolution. From Fig. 4, the depth details can be observed clearly, indicating that the query-based GCPE strategy can contribute to the spatial structure and depth details.
@@
Figure 4: The visualization of heatmap of features and the predicted depth map with/without GCPE.
@@
Figure 5: Overview of SPDecoder (left), which is combined with three spherical geometric priors, and SPAttention (right).
III-C SPDecoder
@@
To balance the global perception and the local details enhancement, we propose the SPDecoder by leveraging spherical priors. As illustrated in Fig. 5, we first send the input features to SPAttention, where we add curve local embedding to provide additional sphere-aware geometric distance. Then, considering the sphere’s continuity, we perform circular rotation to enhance the feature interaction among patches (Fig. 5 ①). We restore the original state through reverse circular rotation (Fig. 5 ②). Afterward, a bipolar re-projection approach based on the spherical equidistortion property is designed to relieve the distortion (Fig. 5 ③). Moreover, circular rotation is operated again to interact with the original features in the vertical direction (Fig. 5 ④). Ultimately, we revert to the original state through reverse operations (Fig. 5 ⑤⑥). We describe these operations in detail as follows.
@@
Figure 6: (a) A projection imaging model with the x-axis (North) as the positive direction. (b) A projection imaging model with the y-axis (North) as the positive direction.
@@
$\left\{\begin{matrix}\lambda^{\prime}=\frac{v^{\prime}}{H}\times\pi\\ \varphi^{\prime}=\frac{u^{\prime}}{W}\times 2\pi\end{matrix}\right.$ (1)
@@
$\left\{\begin{matrix}x=\sin\left(\lambda^{\prime}\right)\\ y=\cos\left(\lambda^{\prime}\right)\\ z=\cos\left(\frac{\pi}{2}-\lambda^{\prime}\right)\cdot\sin\left(\varphi^{\prime}\right)\end{matrix}\right.$ (2)
@@
$\left\{\begin{matrix}\lambda=\arccos\left(\frac{x}{\sqrt{z^{2}+y^{2}}}\right)\\ \varphi=\arcsin\left(\frac{y}{\sqrt{z^{2}+y^{2}}}\right)\end{matrix}\right.$ (3)
@@
$\left(u,v\right)=\left(\frac{\lambda}{\pi}\times H,\frac{\varphi}{2\pi}\times W\right)$ (4)
In particular, we employ the bilinear interpolation function to fill the missing pixels. After BRP operation, the polar regions are projected to the equator, which allows the model to estimate accurate depths with these distortion-free features. Besides, the proposed BRP approach is parameter-free and efficient in transformation, and no additional branches are required.
Circular rotation (CR): The traditional transformer [34] architecture primarily relies on a global self-attention mechanism, which is inefficient when dealing with high-resolution images. To address this, SwinT [30] performs self-attention among divided window-based patches and employs a shifting window operation to facilitate feature interaction between the adjacent patches. Due to the unique seamless property of panoramic images, we implement self-attention interactions between local windows by rotating half the window size in the horizontal direction. Then we rotate back to the original state to align with the initial spatial representation. Since the north and south poles of panoramic images are not adjacent in space, we cannot directly apply CR in the vertical direction. However, benefiting from the BRP module, we can map the features from the vertical to the horizontal direction. Then, through CR, features that are originally in the vertical direction are interacted. Finally, we obtain a feature representation that combines horizontal and vertical directions.
Figure 7: The left image shows an ERP format panoramic image, with each square indicating a segmented patch. Red and blue squares mark patches from the top and bottom with high distortion, while yellow and green squares indicate middle area patches with less distortion. These patches are projected onto the spherical space on the right.
@@
For each patch containing N points, we define $C$ as the CLE within each ${N\times N}$ patch. $\forall P_{1}\left(\lambda_{1},\varphi_{1}\right)$ and $P_{2}\left(\lambda_{2},\varphi_{2}\right)$ (i.e., the position of two points within a patch in spherical coordinates) to calculate the CLE. $c\left(P_{1},P_{2}\right)$ represents the distance between $P_{1}$ and $P_{2}$, where $\lambda\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$/$\varphi\in\left(-\pi,\pi\right)$ represents latitude/longitude, repectively. Through the Haversine formula [35], we can calculate $c\left(P_{1},P_{2}\right)$ as follows:
$h=\sin^{2}\left(\frac{\lambda_{2}-\lambda_{1}}{2}\right)+\cos\left(\lambda_{1}\right)\cdot\cos\left(\lambda_{2}\right)\cdot\sin^{2}\left(\frac{\varphi_{2}-\varphi_{1}}{2}\right)$ (5)
$c\left(P_{1},P_{2}\right)=2R\cdot\arctan\left(\sqrt{\frac{h}{1-h}}\right)$ (6)
@@
TABLE I: Quantitative comparison on the S3D (Structured3D) and Pano3D datasets with the SoTA methods. Blue indicates that our method achieves the best performance.
@@
Bifuse [1] 0.0644 0.0565 0.4099 0.1194 0.9673 0.9892 0.9948
SliceNet [12] 0.1103 0.1273 0.6164 0.1811 0.9012 0.9705 0.9867
Yun et al. [36] 0.0505 0.0499 0.3475 0.1150 0.9700 0.9896 0.9947
S3D [15] PanoFormer [13] 0.0394 0.0346 0.2960 0.1004 0.9781 0.9918 0.9958
EGFormer [2] 0.0342 0.0279 0.2756 0.0932 0.9810 0.9928 0.9964